PREX I Run Plans

From PREX Wiki
Jump to: navigation, search

This was for the PREX-I run in 2010. Run plans for PREX-II and C-REX will be forthcoming.

This was the daily run plans prior to June 21, 2010. Run plans for PREX II are found here:   RunPlanDaily

Daily Run Plan for PREX   (latest at bottom)
Date (2010) Shift ---------- Activity (Responsible) ---------- --- Status ---
03/19 Day pump down; Møller, Compton, septum install; Q2 repair

Insert sieve. (Techs)

done
03/19 Swing Beam restore (B-team) done
03/20 Owl
  • Septum checkout (Michaels: Page when beam imminent & spectrometer ready)

also:

  • Initial look at low current cavity monitors (Ahmed: Page when beam imminent & spectrometer ready)
  • Target alignment (Michaels)
  • VDC checkout (Michaels)
  • Quartz detectors checkout (McNulty, Wexler -- page when VDC tracks seen)
  • GEM checkout (Camsonne -- page when VDC tracks seen, if left dipole is operating)
done

All at 1-2 µA.
Scattered energy ~ beam energy = 1.063 GeV.
Quad settings:

  • Right arm:
    • Q1: dI/dP= 411.2474 Amps/GeV/c
    • Q2: dI/dP= 436.2014 Amps/GeV/c
    • Q3: dI/dP= 433.3897 Amps/GeV/c
  • Left Arm:
    • Q1: dI/dP= 419.8772 Amps/GeV/c
    • Q2: dI/dP= 435.6273 Amps/GeV/c
    • Q3: dI/dP= 434.9682 Amps/GeV/c

"Usual" polarity

03/20 Day Beam off & hall open 9:00 am

Quads repair; Compton install; GEM install; signals debugging. Leave sieves in. (Techs)

done
03/20 Swing
  • Close up hall
  • Benesch tuning (1/2 - 2 hours)
  • Verify quad polarities (Camsonne)
  • Cavity monitors basic function test (Ahmed)
  • Low current cavity monitors setup (Musson, Ahmed)
done
3/21 Owl
  • Low current cavity monitors setup (Musson, Ahmed)
  • Access for GEM installation (Camsonne)
done but GEMs not ready
3/21 Day
  • Spectrometer checkout:
    • Quartz detectors checkout (McNulty, Wexler)
    • Lumi checkout (Mercado) (2 µA on Ta)
    • GEM checkout (Camsonne?)
  • A line tuning (Benesch)
  • Raster checkout
  • Access:
    • Low current cavities
    • Target position
    • Quartz
    • Scintillator alignment
    • LQ2
done
3/21 Swing
  • Low current cavity monitors setup (Michaels, Ahmed)
  • Remove sieves
  • Initial look at sieve out data
  • Replace sieves
done
3/22 Owl
  • Pointing measurement (Riordan)
  • Remove sieves
  • Repeat pointing measurements with no sieves
  • If done or unable to continue before 4 am page Luis Mercado to do lumi checks
  • After 4 am page Dustin McNulty for detector work
done
3/22 Day
  • Optics (Liyanage) Sieve OUT. NB Current limit 2 uA!
done
3/22 Swing
  • Optics continued, sieve IN. NB Current limit 2 uA! (Shift Crew)
done
3/23 Owl
  • Quartz detectors resolution studies (KK, Dustin, John W)
  • Lumi studies (Luis)
done
3/23 Day
  • Remove water target. (Target, Techs)
  • Reconfigure Moller cryo.
  • Beam studies ~ 2pm, study of helicity correlated position differences for spin flipper states. (Dalton, Rupesh, Reza)
water cell removed, rest not finished
3/23 Swing
  • Target pumpdown
done
3/24 Owl
  • Target pumpdown
done
3/24 Day
  • Target cooldown
done
3/24 Swing
  • Tune high current beam into Hall A
done
3/25 Owl
  • Commission lead target, 50 uA (Bob)
  • Lumis and oversampling tests (KK, Luis, Paul)
in progress
3/25 Day
  • Access to fix left dipole
done
3/25 Swing
  • Parity Quality commissioning (Kent, Rupesh, Luis, Dustin)
  • Beam monitor calibrations (Diancheng Wang, Xiaoyan Deng)
postponed
3/26 Owl
  • High Luminosity Detector Commissioning
in progress
3/26 Day
  • Moller Checkout
postponed
3/26 Swing
  • High Current on C12 (thick and thin)
partly done
3/27 Owl no beam
3/27 Day no beam
3/27 Swing
  • High-current on thick C12, adjust HV (Luis or Bob)
  • Finish the commissioning of thick Pb target (Bob, Ahmed)
  • Electronic Commissioning (KK, Luis, Paul, Jon et al)
done, high-current tests continue
3/28 Owl
  • Continue with the above, if necessary.
  • High Luminosity Commissioning (Rupesh, Luis, Jon, Bob)
  • BCM/BPM calibration (Rupesh)
  • Establishing Parity Measurements, Charge feedback, blinding (Rupesh, Jon)
not finished
3/28 Day
  • Comparison Study of Optics Tunes (Seamus, Nilanga, Kiad)
  • Sometime before Monday, submit the ATLIS plan for Dithering test (Rupesh)
done
3/28 Swing
  • Continue Optics if necessary.
  • Continue High Luminosity Commissioning
done
3/29 Owl
  • Mini-production run (shift crew)
  • Stop beam at 07:00 or whenever Ed Folts asks to.
done
3/29 Day
  • Open Hall A to set up Moller and Compton (Ed Folts, Sasha, Sirish, et.al.)
  • Work on GEMs (Alex, Seamus, et.al.)
  • MCC to do beam studies while Hall A open
  • Detector linearity/LED tests (Luis)
  • Setup WAC analysis (Mark, Mindy)
done
3/29 Swing
  • Moller magnet cooldown
  • Optics and acceptance studies (Seamus, Ahmed, Nilanga)
done
3/30 Owl
  • Commission charge feedback (Mark)
  • Current monitor calibration and pedestals determination (Mark)
  • Quasi-production on Empty Target (owl shift crew)
done
3/30 Day
  • Access do work on Moller and Compton.
  • Some GEM installation work. (Alex, Seamus, Kiad, Nilanga)
  • Possible linearity/LED study. (Luis)
done, except linearity
3/30 Swing
  • First Commissioning of Moller -- Beam Checks (Sasha, Javier, Roman, Jay)
done
3/31 Owl
  • Scans of Septum and quad Q1 for tunes A and B. (Bob)
  • Beam Studies of Helicity Correlations
done
3/31 Day
  • Beam Tuning (Jay Benesch)
done
3/31 Swing
  • Either or both of these:
  • a) Moller Commissioning (Sasha, Javier, Roman, Jay)
  • b) Tests of Dithering and other Beam-Related things (Kent, Rupesh)
Moller was done
4/1 Owl
  • If Kent's tests not done on Swing, continue here. (Kent, Rupesh)
  • Beam and Septum scan with new tune (Seamus)
  • Helicity-correlated beam studies (Shift crew)
done
4/1 Day
  • Beam Tuning (Jay)
done, successful
4/1 Swing
  • More beam tuning if necessary
  • Moller Commissioning
done
Friday 2 April Owl
  • Assuming the HRS tune has been frozen
  • Check detector alignment
  • A_T studies
  • Q2 Measurements
partly done, needs more work
Friday 2 April Day
  • Tuning Beam for Compton
done
Friday 2 April Swing
  • 20:00 Obtain sieve runs (optics mini-run plan), Q2 measurements (Bob)
done, but not Q2
Saturday 3 April Owl
  • Obtain sieve runs on Ta (Nilanga mini-run plan)
  • Put in Lead target at low current and do counting-mode runs (Rupesh, Ahmed)
  • Integrate-mode with Lead at 2 or 3 uA (Luis, Jon)
done
Saturday 3 April Day
  • Moller Commissioning (priority activity)
  • Work on getting into production mode (Mark, Rupesh)
done, but not yet in prod.
Saturday 3 April Swing
  • Moller commissioning
done
Sunday 4 April Owl
  • Counting data on tune Y (right arm) super-thin carbon
  • Change to tune B
  • Integrate-mode with Lead at 2 or 3 uA (Luis, Jon)
done
Sunday 4 April Day
  • Integrate-mode with thick carbon 50 uA
  • Compton QWP1 calibration.
done
Sunday 4 April Swing
  • General Compton commissioning and pseudo-production (empty target) parity data at 50 uA.
done
Monday 5 April Owl
  • pseudo-production (empty target) parity data and Compton data at 50 uA.
done
Monday 5 April Day
  • 12:00 Jay to run optics test (~ 30 mins)
  • Moller Spin Dance
  • Set up vertical polarization and measure A_T
done
Monday 5 April Swing
  • Continue to setup up vertical polarization
  • Measurements of A_T on Pb and C12
done
Tuesday 6 April Owl
  • Continue measurements of A_T on Pb and C12
partly done
Tuesday 6 April Day
  • Possible horizontal scan: Tune B with sieve slit (Seamus)
  • When ready: Relocate Quartz Detectors (Dustin, Jack)
postponed: vacuum problems
Tuesday 6 April Swing
  • In restricted access.
  • Any activities that require access.
  • Shift canceled
Wednesday 7 April Owl
  • Shift cancelled
Wednesday 7 April Day
  • Shift cancelled
  • Investigate target vacuum problem
  • Move quartz detectors (Dustin, Jack)
Leak found at O-ring connection between target chamber and collimator box
Wednesday 7 April Swing
  • Shift cancelled
Thursday 8 April Owl
  • Shift cancelled
Thursday 8 April Day
  • Shift cancelled
Target chamber open, found that the 4th foil in Optics target blown off
Thursday 8 April Swing
  • Shift cancelled
Friday 9 April Owl
  • Shift cancelled
Friday 9 April Day
  • To be determined.
  • When ready: Install GEMs
Friday 9 April Swing
  • Cancelled
Saturday 10 April Owl
  • Cancelled
Saturday 10 April Day
  • Cancelled
Saturday 10 April Swing
  • Cancelled
Sunday 11 April Owl
  • Cancelled
Sunday 11 April Day
  • Cancelled
Sunday 11 April Swing
  • Cancelled
Monday 12 April Owl
  • Cancelled
Monday 12 April Day
  • Cancelled
Monday 12 April Swing
  • Cancelled
Tuesday 13 April Owl
  • Cancelled
Tuesday 13 April Day
  • Cancelled
Tuesday 13 April Swing
  • Cancelled
Tungsten collimator modified and installed. Target pump down started and initially successful.
Wednesday 14 April Owl
  • Cancelled
Target pump down.
Wednesday 14 April Day
  • Cancelled -- no shift workers (might need a TO)
  • Target pump down and leak checking.
  • When ready, begin target cooldown. (Dave Meekins and JP)
Done. Vacuum good. Target cold.
Wednesday 14 April Swing
  • Beam restore in Hall A
  • Harp Scan - Shifters
  • Raster size scan (goal is 3x3mm on empty cell) - Shifters.
  • Detector Alignment - Dustin
  • Dithering run with Septum On/Off - Rupesh.
Thursday 15 April Owl
  • Thick Carbon Integrating Mode Run - Rupesh.
  • Thick Carbon Dithering Run - Rupesh + Luis.
  • Take Access: Change to counting mode and put sieve in - Ahmed.
Thursday 15 April Day
  • Beam Studies -- establishing good beam in Hall A and Compton (Jay and B team)
Thursday 15 April Swing
  • Detector Alignment - Dustin.
  • Pb Alignment - Bob.
  • Take access and put sieve in - Seamus.
  • Sieve Alignment - Seamus.
  • Take Access to remove sieve (Seamus) + switch to integrating mode (Jon) - Page Guy and Bob.
  • Heat load test (run up to 500C) - Bob + Guy.

If possible, depending on heat load test:

  • Relative Rate 12C/Pb Runs - Bob / Dustin.
  • Start production (Rupesh, Luis, Kent, Krishna)
Friday 16 April Owl
  • Production
Friday 16 April Day
  • Production
Friday 16 April Swing
  • Production @ 40uA
  • Ramp up to 50uA (Guy + Bob)
Saturday 17 April Owl
  • Production @ 50uA
Saturday 17 April Day
  • Production @ 50uA
  • Moller Checks + Measurement Starting 9am (Sasha)
Saturday 17 April Swing
  • Moller till 7pm (Sasha).
  • Compton (Jay Benesch + Megan).
  • Production @ 50uA.
Sunday 18 April Owl
  • Production @ 50uA
Sunday 18 April Day + Swing
  • Production @ 50uA.
  • Staring 3pm:
  1. Compton test (Jay Benesch + Megan).
  2. Page Dustin + Jon + Bob + Sasha.
  3. Q^2 Measurement (Bob)
  4. Call Compton Experts
  5. Access:
    1. Refill Moller Magnet.
    2. Check Compton HV.
    3. Remove S0.
    4. Go to counting mode (Dustin).
    5. Compton Stuff.
    6. Detector Mapping (Jon).
  6. Detector Alignment (Dustin).
  7. Q^2 Measurement (Bob).
  8. Relative Rate Measurement (Bob).
  9. Access to Return to Integrating Mode.
  • Production @ 50uA
Monday 19 April Owl
  • Production @ 50uA
Monday 19 April Day
  • Production @ 50uA
Monday 19 April Swing
  • Production
Tuesday 20 April Owl
  • Production
Tuesday 20 April Day
  • Production
  • Beam studies from approx. 11 am to 2 pm
  • Hall A Activities:
  • -- Compton 1/4wave stepper (Abdurahim)
  • -- Check R-HRS Q2 and switch to remote mode (Heidi)
  • -- Adjust R-HRS A_T (Dustin)
  • -- Take pedestal data (Jon)
  • -- Check 4th GEM and possibly install 5th (Alex, Jon)
  • -- Test motion of A_T detector R-HRS (Jon)
  • -- Possible Compton shielding / HV check (Megan)
  • -- Ramp Moller magnet, measure stray fields. (Sasha)
  • -- Parasitic work on floor (Yi Qiang)
Tuesday 20 April Swing
  • Tune up Compton
  • Production
Wednesday 21 April Owl
  • Production
Wednesday 21 April Day
  • Production
Wednesday 21 April Swing
  • Production
Thursday 22 April Owl
  • Production
Thursday 22 April Day
  • Beam studies. Transverse Phase Space Study in Wien. (Reza, Mark, Rupesh)
  • At ~2 pm start Moller commissioning
Thursday 22 April Swing
  • Finish Moller (10 hours)
  • Resume Production
  • Don't forget the beam recovery procedure
Friday 23 April Owl
  • Production
Friday 23 April Day
  • Q2 Studies (3 hrs) -- Dustin, Bob
  • GEM studies are only parasitic.
Friday 23 April Swing
  • Production
Saturday 24 April Owl
  • Production on thick C12
Saturday 24 April Day
  • Production on thick C12
  • At about 2 pm switch to 120 Hz, still C12
  • Run briefly on C12 at 120 Hz
  • Switch to thick Pb target (120 Hz)
Saturday 24 April Swing
  • Production on Pb at 120 Hz
Sunday 25 April Owl
  • Production
Sunday 25 April Day
  • Q2 measurements - Bob, Dustin
  • Otherwise production
done
Sunday 25 April Swing
  • Production
Monday 26 April Owl
  • Production
Monday 26 April Day
  • Wien flip
Wien flip postponed until Tues
Monday 26 April Swing
  • Production
Tuesday 27 April Owl
  • Production
Tuesday 27 April Day
  • Beam Studies
  • Wien Flip
  • Checks of laser at polarized source (Kent and EGG)
  • Adiabatic Damping Study (Yves Roblin)

Done

Tuesday 27 April Swing
  • Compton Tune
  • 4 hrs on Thick Carbon (RD=0, IT=500us)
  • flip half-wave plate
  • 4 hrs on Thick Carbon
Wednesday 28 April Owl
  • Finish Swing Plan
  • Yves Roblin Measurement (if possible)
  • Run on Carbon (RD=0, IT=1535)
Wednesday 28 April Day
  • Moller
Wednesday 28 April Swing
  • Finish Moller
  • Production
Thursday 29 April Owl
  • Production
Thursday 29 April Day
  • Production
Access to repair target leak
Thursday 29 April Swing
  • Production
Could not close hall due to faulty fire suppression system
Friday 30 April Owl
  • Production
Could not close hall due to faulty fire suppression system
Friday 30 April Day
  • 08:00 Beam Stops for Open House
Friday 30 April Swing
  • No Beam
  • Injector studies
Friday 30 April Owl
  • No Beam
  • Injector studies
Saturday 1 May Day
  • Open House
  • Detector work
  • Compton work
Saturday 1 May Swing
  • Restore Beam.
  • Check Beam Properties. (shift)
  • Pulsed Beam Test. (Mark, Luis)
Sunday 2 May Owl
  • Continue with Above Tests
  • Raster and Pb Target Tests (Bob)
Sunday 2 May Day
  • Continue Above Tests if necessary
  • Start Production
  • Get cavities to work at high current. (Ahmed)
  • (The check of cavities at low current is postponed.)
Sunday 2 May Swing
  • Production
Monday 3 May Owl
  • Production
Monday 3 May Day
  • Hall B Pass change 3 -> 5 until ~1:00PM
  • Compton vacuum, LHRS dipole UPS replacement during the down
  • Bake compton chicane overnight during production, run beam straight through
  • Return to PbD #2 production
  • Large asymmetry run (Jon)
  • Baking started, HRS work completed by Heidi
Monday 3 May Swing
  • Production
Tuesday 4 May Owl
  • Production
Tuesday 4 May Day
  • Work on switch for Pockels cell (start 9:00AM)
  • Complete Compton beam line bake-out, leak check, cavity alignment
  • Pockels cell switch work complete
  • Compton vacuum reopened
Tuesday 4 May Swing
  • Production
  • Continued Compton realignment/laser polarization measurement
Wednesday 5 May Owl
  • Production
  • Stopped work on Compton - still not complete
  • Changed beam spot size to 200x200um
Wednesday 5 May Day
  • 12 hours Moller tests, start 9:00AM (Sasha, Jixie)
Wednesday 5 May Swing
  • Finish Moller tests, 21:00
  • Raster clock tests 22:00-24:00 (KK, Bob, Jon)
  • Production as per raster clock tests
Thursday 6 May Owl
  • Production
Thursday 6 May Day
  • Beam Studies (8 hours controlled access starting 08:15)
    • Finish Compton cavity work (Sirish, Abdurahim)
    • Start Compton pumping, replace leaky upstream valve? (John Heckman et al)
    • Remove GEM(s)?? (Alex, Kiad)
    • Compton Boron installation (Megan)
  • Continued compton cavity alignment
  • Boron installed
  • INFN GEMs removed to EEL, (additional UVA GEM installed?
Thursday 6 May Swing
  • Raster clock tests
  • Production
  • CHL trip
Friday 7 May Owl
  • Raster clock tests (cont?)
  • Production
  • Tests aborted due to network switch failure in the hall
Friday 7 May Day
  • Production
  • Wien flip, QE scan, spot move (nominally 9:00AM)
  • Wien flip, QE scan, spot move completed.
  • Compton valve work continued until 2:00
  • Started detector alignment checks
  • Q2 and establish low-current cavities (Dustin)
Friday 7 May Swing
  • Tests by KK et al
  • Production
Saturday 8 May Owl
  • Production
Saturday 8 May Day
  • Production
  • 2:00PM - Turn off cavity baking
  • 3:00PM - Leak check, turn on ion pumps
  • 4:00PM - Routine Moller Measurement
  • Compton vacuum reestablished
  • Replaced valve opens with sufficient vacuum levels
Saturday 8 May Swing
  • Moller
  • Moller measurement gives 91% polarization
Sunday 9 May Owl
  • Production
Sunday 9 May Day
  • Compton ready?
  • Production
  • Cavity locks at 3kW!
  • Valve upstream of chicane will not open
Sunday 9 May Swing
  • Production
Monday 10 May Owl
  • Production
  • Left dipole down
  • Compton valve stuck
Monday 10 May Day
  • Fix left dipole
  • Fix Compton valve
  • Fix target vacuum
  • Left dipole down
  • Compton valve fixed and tested
  • Target vacuum bad
Monday 10 May Swing
  • Standby
  • Empty or C target oversampling runs if valves can be opened
  • 1-arm production if target vacuum < 5E-5
  • Left dipole down
  • Target vacuum bad
Tuesday 11 May Owl
  • Standby
  • Empty or C target oversampling runs if valves can be opened
  • 1-arm production if target vacuum < 5E-5
Tuesday 11 May Day
  • Wien flip
  • Beam studies -- injector studies
  • Raster setup
  • Fix left dipole
  • Fix target vacuum
Tuesday 11 May Swing
  • Compton tune
  • Oversampling tests
  • Left dipole down
  • Right dipole NMR down
  • Target vacuum bad
Wednesday 12 May Owl
  • Production
  • Left dipole down
  • Right dipole NMR down
  • Target vacuum bad
Wednesday 12 May Day
  • Repair left dipole PS
  • Repair right dipole NMR
  • Repair right Q3 bad bit
  • Target vacuum bad
Wednesday 12 May Swing
  • Standby
  • Left dipole down
  • Left Q3 controls problem
  • Target vacuum bad
Thursday 13 May Owl
  • Standby
  • Left dipole down
  • Left Q3 controls problem
  • Target vacuum bad
Thursday 13 May Day
  • Repair left dipole PS
  • Repair left Q3 controls
  • Find and stop target vacuum leak
  • Warm up target
Thursday 13 May Swing
  • Warm up target
Friday 14 May Owl
  • Pump down target
Friday 14 May Day
  • Pump down target
Friday 14 May Swing
  • Cool target
  • Re-establish beam
  • Raster tests
  • Target vacuum good
  • Both spectrometers good
Saturday 15 May Owl
  • Production
Saturday 15 May Day
  • Access for left dipole regulation
  • Left dipole regulation failed
  • Target vacuum increase, then decrease with added heating
Saturday 15 May Swing
  • Access for left dipole regulation
  • Re-establish beam
  • Møller
  • Left dipole regulation bad
  • Target vacuum varying, mostly < 1e5
Sunday 16 May Owl
  • Møller
  • Left dipole regulation bad
  • Target vacuum varying, mostly < 1e5
Sunday 16 May Day
  • Access for left dipole regulation
  • Target vacuum varying, mostly < 1e5
Sunday 16 May Swing
  • Production
  • Background test (L PMTs on)
  • Left dipole regulation bad
  • Target vacuum varying, mostly < 1e5
Monday 17 May Owl
  • Production
Monday 17 May Day
  • Controlled Access for Left Dipole fix
  • Moller refilled
  • Left dipole working for now
  • Suspect bad silicon-controlled rectifier (SCR)
  • There are 12 needed and only 6 spares
  • Looking for spares and will work during beam studies Tue.
Monday 17 May Swing
  • Production with both arms!
Tuesday 18 May Owl
  • Production (hopefully with both arms!)
  • No Left Dipole issues all night
Tuesday 18 May Day
  • Beam Studies, beginning 9am (8hrs)
  • Wien Flip
  • Maintenance on the Left Dipole Power supply
  • Wien Flip completed before beam studies
  • Beam studies completed
  • Measurements performed on Left Dipole Power supply
Tuesday 18 May Swing
  • Production
  • Oversample tests (Rupesh)
  • Compton not done
  • Oversample tests not done
Wednesday 19 May Owl
  • Production
  • Oversample tests
  • Oversample tests completed
Wednesday 19 May Day
  • Compton Tuning (1-2hrs, start around 10am)
  • Raster tests during 'target vacuum recovery'
  • Production
  • Injector tuning did not help Compton rates
  • Jay and Yan are asking for advice
  • Compton tuning during beam studies tomorrow
Wednesday 19 May Swing
  • Production
  • Q2 and detector alignment (2-3hrs, around 8pm)
  • short Q2 measurement
  • Target vacuum tripped valves closed -> no beam
Thursday 20 May Owl
  • Production
  • No beam - target operator only
Thursday 20 May Day
  • Beam Studies (start at 8:30am, afternoon for Compton tuning)
  • Some measurements on left dipole power supply by Jack in Hall
  • Production
  • Compton tuning needs to be rescheduled
  • Jack did some work in hall
Thursday 20 May Swing
  • Canceled
We lost the target vacuum. Will need to warm up target and replace O-ring after letting radiation subside. Plan to replace O-ring on Monday May 24.
Friday 21 May Owl
  • Canceled
Friday 21 May Day
  • Canceled
Friday 21 May Swing
  • Canceled
Saturday 22 May Owl
  • Canceled
Saturday 22 May Day
  • Canceled
Saturday 22 May Swing
  • Canceled
Sunday 23 May Owl
  • Canceled
Sunday 23 May Day
  • Canceled
Sunday 23 May Swing
  • Canceled
Monday 24 May Owl
  • Canceled
Monday 24 May Day
  • Various tests
    • Pedestal runs
    • linearity studies
    • detector time constant tests
    • GEM studies
  • Target repairs - replace o-ring with silicon one temporarily
  • Keep target warm and pump down
  • Leak check
Monday 24 May Swing
  • Pedestal runs
Tuesday 25 May Owl
  • Canceled
Tuesday 25 May Day
  • Beam Studies (until between 10am and noon)
  • Pedestal runs
Kalres O-ring installed ahead of schedule. Pumping down.
Tuesday 25 May Swing
  • Pedestal runs
Wednesday 26 May Owl
  • Canceled
Wednesday 26 May Day
  • Accelerator Vacuum Group in Hall for Leak checks
  • Switch box arrived for activate remotely the heating and cooling of thetarget
  • Found a leak, but pressure stable at 675 mtor
  • Electrician will install the appropriate outlet; R. Michaels is working on the remote control.
Wednesday 26 May Swing
  • Canceled
Thursday 27 May Owl
  • Canceled
Thursday 27 May Day
  • Accelerator Vacuum Group in Hall for Leak checks
Thursday 27 May Swing
  • Canceled
Friday 28 May Owl
  • Canceled
Friday 28 May Day
  • Shift workers on Standby
  • Accelerator Vacuum Group in Hall for Leak checks
Friday 28 May Swing
  • Canceled
Saturday 29 May Owl
  • Canceled
Saturday 29 May Day
  • Canceled
Saturday 29 May Swing
  • Canceled
Sunday 30 May Owl
  • Canceled
Sunday 30 May Day
  • Canceled
Sunday 30 May Swing
  • Canceled
Monday 31 May Owl
  • Canceled
Monday 31 May Day
  • Canceled
Monday 31 May Swing
  • Canceled
Tuesday 1 June Owl
  • Canceled
Tuesday 1 June Day
  • Canceled
Tuesday 1 June Swing
  • Canceled
Wednesday 2 June Owl
  • Canceled
Wednesday 2 June Day
  • Canceled
Wednesday 2 June Swing
  • canceled
Thursday 3 June Owl
  • Canceled
Thursday 3 June Day
  • Cooling down target (more likely to start on Swing)
  • Day shift on standby.
Thursday 3 June Swing
  • Swing shift workers are needed.
    • PREX detector in counting mode
    • resume beam in Compton chicane ( MCC )
    • harp scan for beam profile and beam position ( Shift crew - MCC )
    • raster check ( Shift crew )
    • Long down target recovery procedure ( Shift crew )
  • Alignment checks, Q2 ( Dustin Mc Nulty ).
    • put sieve in
    • take alignement data
    • put sieve out
    • Q2 measurement
    • Ask for access to go to integrating mode ( Jon Wexler or Amhed Zafar ) and remove Sieve
  • Restart Production
  • Production
Friday 4 June Owl
  • Production
Friday 4 June Day
  • Production
Friday 4 June Swing
  • Production
Saturday 5 June Owl
  • Production
Saturday 5 June Day
  • Production
Saturday 5 June Swing
  • Production
Sunday 6 June Owl
  • Production
Sunday 6 June Day
  • Production
Sunday 6 June Swing
  • Production
Monday 7 June Owl
  • Production
Monday 7 June Day
  • Moller Measurement (noon to 4pm)
  • Big Change asymmetry run
  • Production
Monday 7 June Swing
  • Production
Tuesday 8 June Owl
  • Production
Tuesday 8 June Day
  • Production
  • Wien Flip (9 AM)
  • Beam study (9-11 AM)
Tuesday 8 June Swing
  • Production
Wednesday 9 June Owl
  • Production
Wednesday 9 June Day
  • Production
Wednesday 9 June Swing
  • Production
Thursday 10 June Owl
  • Production
Thursday 10 June Day
  • Production
Thursday 10 June Swing
  • Production
Friday 11 June Owl
  • Production
Friday 11 June Day
  • Production
  • Wien flip
Friday 11 June Swing
  • Production
Saturday 12 June Owl
  • High Statistics Lead Inelastics, Qsq|GEM test, Calibration of the cavities with the amplifier
all done
Saturday 12 June Day
  • Backgrounds: Dipole Scan, Transverse Polarization Measurements.
Finished left dipole scan, right dipole scan is pending
Saturday 12 June Swing
  • Transverse Polarization Measurements. Carbon
Sunday 13 June Owl
  • Transverse Polarization Measurements. Lead
Sunday 13 June Day
  • Transverse Polarization Measurements. Lead

4494 32k (beam current stepping up, might not be good) 4495 78k 4496 73k 4497

Already took 580K multiplets, 60K of them may not be good data
Sunday 13 June Swing
  • If right Q2 fixed, take 100K more A_T data then go back to production. Otherwise go back to single arm production
Monday 14 June Owl
  • Continue Transverse Polarization Measurements. Blinded tube test
Monday 14 June Day
  • Production

Remaining auxiliary measurements

  • PREx det. current calibrations (Monday swing)
  • Finish dipole scan (Tuesday owl)
  • Oversampling by 2 at 120 Hz (put off/cancelled)
  • Beam induced linearity studies (put off/cancelled)
  • Q2 measurements (relative using T5 during dipole scan)
  • MOLLER commissioning (if can't do anything else - need repair? - see Jixie or Roman)
Monday 14 June Swing
  • Production
  • PREx det. current calibrations (Partially done)
  • large charge asymmetry run couldn't be done b/c MCC couldn't set RHWP to 2500 (see Rupesh' log entry)
Tuesday 15 June Owl
  • Production
  • Finish dipole scan (Done)
  • Modified dipole scan completed overnight (see Bob's entries 1 and 2)
Tuesday 15 June Day
  • Beam Studies

expected to deliver beam to Hall A - NO AUTOMATIC ACCESS

  • Production
  • large charge asymmetry run taken (see run 4574)
  • Beam was lost during beam studies after all
    • Megan and Alex went in to check on the Compton
    • Jixie and Roman did some work on the Moller, including a refill
    • Jack and Howard worked on the RHRS dipole which had issues
Tuesday 15 June Swing
  • Production
Wednesday 16 June Owl
  • Production
  • went to single arm running after right dipole power supply trip
Wednesday 16 June Day
  • Right arm dipole repairs
  • Put silicone on target
  • Production
Wednesday 16 June Swing
  • Begin with PbD#3 - check temperatures
    • If too high - switch back to PbD#1
  • Production - single arm running
  • Beam not back until ~11pm
  • Try to do beam recovery before owl starts
Thursday 17 June Owl
  • Production (single arm)
Thursday 17 June Day
  • Right dipole repairs
  • Wien Flip (10am)
  • Wien Flip done this morning
  • Right dipole working for now
  • Target warm up flashed off N2 ice - target cooling down
Thursday 17 June Swing
  • Trouble sending 70 uA beam
  • BCMs needed calibration
Friday 18 June Owl
  • Production
  • 70 uA beam transport established
  • Problem with RHRS rates discovered (25x smaller than LHRS)
Friday 18 June Day
  • Production
  • End of Run Party 2-4pm in the ResFac Great Room
  • RHRS sieve discovered to be partially in; Howard went in and moved it out
Friday 18 June Swing
  • Production
Saturday 19 June Owl
  • Production
Saturday 19 June Day
  • Production
Saturday 19 June Swing
  • Production
Sunday 20 June Owl
  • Production
Sunday 20 June Day
  • Production
Sunday 20 June Swing
  • Production
Monday 21 June Owl
  • Production
Monday 21 June Day
  • Dismantle detectors - see Dustin or Alex
Tuesday 21 June Day
  • Counting house cleaning party beginning at 8am!